OpenStudy (anonymous):

Let f be the function given by f(x) = \[\sqrt{x^4-16x^2}\] ---- find the slope of the line normal to the graph of f at x=5

OpenStudy (mertsj):

The first derivative is:\[\frac{2x(x^2-8)}{\sqrt{x^2(x^2-16)}}\]

OpenStudy (mertsj):

That is the slope function. Replace x with 5 to find the slope of the tangent at x = 5

OpenStudy (anonymous):

oh so i just plug it in and that gives me the slope?

OpenStudy (mertsj):

\[\frac{2(5)(25-8)}{\sqrt{25(25-16)}}=\frac{170}{225}=\frac{34}{45}\]

OpenStudy (mertsj):

That is the slope of the tangent. The slope of the perpendicular is: \[\frac{-45}{34}\]

OpenStudy (anonymous):

thank you!

OpenStudy (mertsj):

yw